OpenStudy (anonymous):

suppose xy=3 and dy/dt=4. find dx/dt when x=-2

OpenStudy (anonymous):

I got 16/3

OpenStudy (anonymous):

(dy/dt)/(dx/dt)=dy/dx Solve for y when x =-2 for starters then find dy/dx implicitly.

OpenStudy (anonymous):

I don't know how to do that? :/

OpenStudy (anonymous):

Can you tell me how like generally so i can solve it?/

OpenStudy (anonymous):

Starting by solving for y?

OpenStudy (anonymous):

doesn't y=3?

OpenStudy (anonymous):

xy=y?

OpenStudy (anonymous):

when x=2?\[xy=3\]\[-2y=3\]

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

yes. Now you have an ordered pair (-2, -3/2) which will come in handy later.

OpenStudy (anonymous):

Can you differentiate implicitly?

OpenStudy (anonymous):

No, that's what I don't understand at all.

OpenStudy (anonymous):

Well, you could say that all differentiation is implicit, because when we differentiate implicitly we treat the y as a variable to be differentiated, with all the usual rules in effect. \[\frac{ d }{ dx }(y)=\frac{ dy }{ dx }\] What we'll do, though instead of using the product rule on xy, is solve for y and differentiate. That should be easier. So,\[xy=3\]\[y=\frac{ 3 }{ x }\]Can you differentiate that?

OpenStudy (anonymous):

You could also write the x with a negative exponent if you prefer the power rule

OpenStudy (anonymous):

-3/x^2?

OpenStudy (anonymous):

Yeah. When you substitute x=-2, you get the slope, or dy/dx, which I said earlier was equal to (dy/dt)/(dx/dt)

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

4=3/4*dx/dt?

OpenStudy (anonymous):

oh -3/4

OpenStudy (anonymous):

otherwise you're right. Just solve.

OpenStudy (anonymous):

19/4

OpenStudy (anonymous):

You dropped the negative sign, and where did you get 19??

OpenStudy (anonymous):

4+3/4?

OpenStudy (anonymous):

my calculator gave me 19/4

OpenStudy (anonymous):

You weren't supposed to add. 4=dy/dt. -3/4=dy/dx. dy/dx=(dy/dt)/(dx/dt). dx/dt=?

OpenStudy (anonymous):

-16/3, I got it...